# Causal Mechanisms of Subpolar Gyre Variability in CMIP6 Models

### v2

@@ 4.2 Causal Effect
To compute the causal effect for the links we need to set the lags at which they are considered. If more than five models indicate a link at a certain lag to be significant, it is included for the computation of the causal effect. This means that for all theoretical links between variables we take into account both direct and lag-1 links, except for SPGSSS where only links at lag-1 are considered. Furthermore, links from SSS, MLD and Rho to themselves at lag-1 are used, while for SubT lag-2 is considered in addition. For SPG lags up to five years are taken into account.
@@
In Figure 5 the causal effect of each link is shown for models with and without that link being significant following the causal discovery step (values for individual models are given in the SI). For the direct SSSMLD link (orange violins) all (significant) models identify the correct sign, with an increase in salinity leading to a deeper mixed layer. This is in line with the results found for the mechanisms of convection found in Section 3. Also for the direct MLDSubT link (green violins) most significant models identify the right direction with an increase in MLD leading to a cooling of SubT. At lag-1 both these links are found with the opposite sign in most models where they are significant, contrary to the theoretical model. This sign-change is only found when taking into account the longer memory (beyond lag-1) in SPG and SubT, meaning the absence of it in the theoretical model is likely due to the use of correlation to establish the links (Born and Mignot, 2012). Physically these links could be related to the atmospheric dampening of the surface signal, which is not included.
@@
For the proposed SubTRho link (red violins) models disagree on the sign and value of the contemporaneous link, with a large spread for the significant models. We find a positive link at lag-1, whereas in the theoretical model the link is negative, with a cooling of SubT leading to an increase of Rho. This likely is due to the different effect of mixing on the surface density versus the density at depth. Where the mixing of the water column leads to an increase of density at depth, it actually reduces the density at the surface corresponding to the positive link found at lag-1.
@@
The sign of the SPGSSS link (blue violin) is very consistent in the models where it is found to be significant (but not overall). For those models it shows a negative feedback with an increase in SPG leading to a reduction in SSS a year later. The RhoSPG link (purple violins) is less consistent, but most significant models indicate that a reduction in Rho leads to a strengthening of SPG, either direct or at lag-1. This is not in agreement with the theoretical mechanism, where an increase in density strengthens the gyre circulation. It indicates that the models do not capture the feedback of the density in the gyre centre to its strength. The RhoSPG link is based on thermal wind balance (Born and Stocker, 2014), with an increase in the density difference between edge and centre of the gyre increasing the strength of the circulation. This absence could be related to the use of the barotropic streamfunction as a measure of the gyre strength, which does not fully capture the baroclinic nature of the link. It also hints at there being other processes in CMIP6 models, such as the atmospheric circulation, which act as confounding factors.
@@
Since both links through Rho are identified with the opposite sign of the theoretical model, it is valuable to study the system when leaving out Rho, i.e. the hypothesis of a causal link from SubT to SPG. Causally the hypothesis omitting Rho is equivalent to the studied system (Pearl et al., 2016), although the effect on lag times is hard to predict. Studying this system (see SI) shows large disagreement between models on the sign of the SubTSPG link at lags of one and two years. At lag-1 slightly more significant models indicate a positive link, i.e. a cooling of SubT leading to a weakening of the SPG, which is the opposite effect of the theoretical mechanism. However, at lag-2 slightly more models show a negative effect, in line with the hypothesis. Thus, the models do not agree on the SubTSPG link, whether taking into account the step through Rho or not.
@@
There is one model, CESM2, for which each link is significant either direct or at lag-1. For the SSSMLDSubT links CESM2 indicates significant contemporaneous links, whereas for SubTRhoSPG the significant links are found at lag-1. The sign of the links to and from Rho both have the opposite sign to the theoretical mechanism, which retains the hypothesized negative feedback loop. When omitting Rho, CESM2 has a significant negative SubTSPG link at lag-2, being in line with the negative feedback loop. Interestingly, CESM2 is one of the models in which Swingedouw et al. (2021) found an abrupt change of temperature in the subpolar gyre region linked to a collapse of convection. This collapse of convection has more recently also been found in CESM2 large ensemble runs (Gu et al., 2024). This is evidence that this model can exhibit bistability of the subpolar gyre, where the identification of links that are (up to a certain level) in line with the proposed mechanism increases confidence in the model.
@@ 5 Conclusions
We have studied the mechanisms of SPG tipping behaviour in CMIP6 models. We started with considering the mechanisms of convection (A) (Figure 1), for which we found that, in line with observations, either an increase of salinity or a decrease of temperature causes convection (mechanism (A1)), i.e. an increase in mixed layer depth. Here, the effect of salinity is found to be slightly stronger than that of temperature. In contrast to the clear representation of these causes of convection, the feedback of mixed layer depth to the sea surface temperature is not identified robustly in CMIP6 models (mechanism (A2)). This is due to the lack of causal sufficiency of the tested hypothesis, with common drivers of temperature and salinity not being included. There are several candidates for common drivers, for example the ocean circulation (e.g. Kim et al., 2021) or the state of the atmosphere (e.g. Khatri et al., 2022), where these in turn also interact with each other. Therefore, we turned to testing mechanism (B), which is causally sufficient. This mechanism, proposed by Born and Mignot (2012), can describe bistability, and thus tipping, of subpolar gyre convection. It includes a feedback between salinity and the gyre circulation, which is in part due to eddies.
@@
We verified whether the links between the variables in mechanism (B) are well represented in CMIP6 models. The proposed mechanism, shown in Figure 1, contains a negative feedback loop on short timescales and a positive one on longer timescales. The interaction between salinity, mixed layer depth and the subsurface temperature is relatively well captured in many CMIP6 models. This shows that convection is indeed driven by surface density and is in line with the results for the convection mechanism (A1). The feedbacks between the subsurface temperature, density and strength of the subpolar gyre are present in fewer models, often with conflicting signs between models. The positive link from the temperature at depth to density at the surface can physically be explained by the upward mixing of lower density water. The mostly negative link from density to the strength of the gyre is not in line with the theoretical mechanism, indicating that models do not contain the expected interaction between density and the gyre circulation.
@@
The lagged feedback from the gyre circulation to salinity is not found in any of the models and as a consequence the delayed positive feedback loop is missing. The long memory of the circulation strength can (partly) explain this, as lagged regression (used in Born and Mignot (2012)) can lead to identifying too many significant relationships (McGraw and Barnes, 2018). At a lag of two years we find a weakening effect of the circulation strength on itself in most significant models (shown in SI), which would act to increase the salinity on a timescale of around 3 years, i.e. a positive feedback loop. This is a shorter timescale than identified by Born and Mignot (2012), which can be due to the gyre-memory biasing their analysis. The gyre-memory can be explained by the transport of waters from outside the Labrador sea, which retains similar properties for multiple years, as seen during the Great Salinity Anomalies (Gelderloos et al., 2012). In CESM2 this lagged link of the gyre strength is found to be significant, providing evidence for both a negative and delayed positive feedback loop being present in this model. There might be other models in which such a mechanism is present, but where noise reduces the significance of one or more links.
@@
The link that contradicts most with the bistability mechanism (B) is that from density to the gyre strength. This is one of the links that relies heavily on a good parametrization of heat and salt transport by ocean eddies, and hence it is not surprising that this interaction is not identified clearly. Even though eddies are better resolved in high-resolution models, we did not find a discernible effect of model resolution on the identification and strength of this link. Another partial explanation is that the barotropic streamfunction, which is used as a measure of the strength of the subpolar gyre, is not the most suitable variable to use for the interaction. The streamfunction is computed by vertical integration, which removes baroclinic effects, whereas it is exactly these baroclinic effects that are relevant for the positive feedback loop (Born and Stocker, 2014). Together, these likely explain the absence of this interaction in most CMIP6 models.
The absence of the interaction between density and gyre strength means that we cannot be certain of the existence of the bistability mechanism in many CMIP6 models, and thus whether they are able to exhibit tipping behaviour in SPG convection. Nevertheless, several models have shown abrupt shifts in temperature in the subpolar grye region, sometimes accompanied by a shutdown of convection. CESM2, in which Gu et al. (2024) have identified such abrupt shifts in temperature and convection, is the only model where all links in some form were found to be present, indicating that these abrupt shifts may indeed be linked to the proposed bistability mechanism (Born and Mignot, 2012). For other models that exhibited shifts in temperature, we did not find all links to be significant, meaning they could also be a nonlinear response to forcing instead of tipping behaviour.

### v3

@@ 4.2 Causal Effect
To compute the causal effect for the links we need to set the lags at which they are considered. If more than five models indicate a link at a certain lag to be significant, it is included for the computation of the causal effect. This means that for all theoretical links between variables we take into account both instantaneous and lag-1 links, except for SPGSSS where only links at lag-1 are considered. Furthermore, links from SSS, MLD and Rho to themselves at lag-1 are used, while for SubT lag-2 is considered in addition. For SPG lags up to five years are taken into account.
@@
In Figure 5 the causal effect of each link is shown for models with and without that link being significant following the causal discovery step (values for individual models are given in Fig. S5). Distinguishing between significant and non-significant models allows to separate the models in which we do expect to find the relevant dynamics from those where we do not (e.g. because they show little convection in the Labrador Sea). For the instantaneous SSSMLD link (orange violins) all (significant) models identify the correct sign, with an increase in salinity leading to a deeper mixed layer. This is in line with the results found for the mechanisms of convection found in Section 3. Also for the instantaneous MLDSubT link (green violins) most significant models identify the right direction with an increase in MLD leading to a cooling of SubT. At lag-1 both the SSSMLD and MLDSubT link are found with the opposite sign in most models where they are significant, contrary to the theoretical model. This sign-change is only found when taking into account the longer memory (beyond lag-1) in SPG and SubT, meaning the absence of it in the theoretical model is likely due to the use of correlation to establish the links (Born and Mignot, 2012). Physically these links could be related to the atmospheric dampening of the surface signal, which is not included.
@@
For the proposed SubTRho link (red violins) models disagree on the sign and value of the instantaneous link, with a large spread for the significant models. We find a positive link at lag-1, whereas in the theoretical model the link is negative, with a cooling of SubT leading to an increase of Rho. This likely is due to the different effect of mixing on the surface density versus the density at depth. Where the mixing of the water column leads to an increase of density at depth, it actually reduces the density at the surface corresponding to the positive link found at lag-1.
@@
The sign of the SPGSSS link (blue violin) is very consistent in the models where it is found to be significant (but not overall). For those models it shows a negative feedback with an increase in SPG leading to a reduction in SSS a year later. The RhoSPG link (purple violins) is less consistent, but most significant models indicate that a reduction in Rho leads to a strengthening of SPG, either instantaneous or at lag-1. This is not in agreement with the theoretical mechanism, where an increase in density strengthens the gyre circulation. It indicates that the models do not capture the feedback of the density in the gyre centre to its strength. The RhoSPG link is based on thermal wind balance (Born and Stocker, 2014), with an increase in the density difference between edge and centre of the gyre increasing the strength of the circulation. This absence could be related to the use of the barotropic streamfunction as a measure of the gyre strength, which does not fully capture the baroclinic nature of the link. It also hints at there being other processes in CMIP6 models, such as the atmospheric circulation, which act as confounding factors.
@@
Since both links through Rho are identified with the opposite sign of the theoretical model, it is valuable to study the system when leaving out Rho, i.e. the hypothesis of a causal link from SubT to SPG. Causally the hypothesis omitting Rho is equivalent to the studied system (Pearl et al., 2016), although the effect on lag times is hard to predict. Studying this system (see Figs. S7-S9) shows large disagreement between models on the sign of the SubTSPG link at lags of one and two years. At lag-1 slightly more significant models indicate a positive link, i.e. a cooling of SubT leading to a weakening of the SPG, which is the opposite effect of the theoretical mechanism. However, at lag-2 slightly more models show a negative effect, in line with the hypothesis. Thus, the models do not agree on the SubTSPG link, whether taking into account the step through Rho or not.
@@
There is one model, CESM2, for which each link is significant either instantaneous or at lag-1. For the SSSMLDSubT links CESM2 indicates significant instantaneous links, whereas for SubTRhoSPG the significant links are found at lag-1. The sign of the links to and from Rho both have the opposite sign to the theoretical mechanism, which retains the hypothesized negative feedback loop. When omitting Rho, CESM2 has a significant negative SubTSPG link at lag-2, being in line with the negative feedback loop.
If we do not take into account the significance of the links, but only consider the causal effect, four other models have the same sign as CESM2 for all links; CESM-FV2, CESM2-WACCM, MRI-ESM2-0 and NorESM2-LM. The models in which Swingedouw et al. (2021) found abrupt temperature shifts in the SPG region are CESM2-WACCM, MRI-ESM2-0 and NorESM2-LM. This identifies a clear connection between the presence of mechanism (B) and the possibility for abrupt shifts in CMIP6 models. That not all links in these models are found to be significant is likely linked to internal variability obscuring the signal. The key links which set this set of models apart are those to and from the SPG strength. When constraining on these three links only, i.e. negative SPGSSS and RhoSPG positive at lag-0 and negative at lag-1, we find this set of five models together with ACCESS-CM2. In this last model the SubTRho link at lag-1 is negative instead of positive, and it is unclear what this means for the possibility of an abrupt shift in the SPG as this model was not considered in Swingedouw et al. (2021).
@@ 5 Conclusions
We have studied the mechanisms of SPG tipping behaviour in CMIP6 models. We started with considering the local mechanisms of convection (A) (Figure 1), for which we found that, in line with observations, either an increase of salinity or a decrease of temperature causes convection (mechanism (A1)), i.e. an increase in mixed layer depth. Here, the effect of salinity is found to be slightly stronger than that of temperature. In contrast to the clear representation of these causes of convection, the feedback of mixed layer depth to the sea surface temperature is not identified correctly in CMIP6 models (mechanism (A2)). This is due to the lack of causal sufficiency of the tested hypothesis, with common drivers of temperature and salinity not being included. There are several candidates for common drivers, for example the ocean circulation (e.g. Kim et al., 2021) or the state of the atmosphere (e.g. Khatri et al., 2022), where these in turn also interact with each other. Therefore, we turned to testing mechanism (B), which is causally sufficient. This mechanism, proposed by Born and Mignot (2012), can describe bistability, and thus tipping, of subpolar gyre convection. It includes a feedback between salinity and the gyre circulation, which is in part due to eddies. Instead of focusing on the location where individual models show convection, in this part of the analysis we consider only the Labrador Sea.
@@
We verified whether the links between the variables in mechanism (B) are well represented in CMIP6 models. The proposed mechanism, shown in Figure 1, contains a negative feedback loop on short timescales and a positive one on longer timescales. The interaction between salinity, mixed layer depth and the subsurface temperature is relatively well captured in many CMIP6 models. This shows that convection is indeed driven by surface density and is in line with the results for the local convection mechanism (A1). The feedbacks between the subsurface temperature, density and strength of the subpolar gyre are present in fewer models, often with conflicting signs between models. The positive link from the temperature at depth to density at the surface can physically be explained by the upward mixing of lower density water. The mostly negative link from density to the strength of the gyre is not in line with the theoretical mechanism, indicating that models do not contain the expected interaction between density and the gyre circulation. One reason for not identifying all links is due to some models not showing (deep) convection in the Labrador Sea, in particular the CanESM5 and EC-Earth models. However, even for models with a link being significant we do not identify all links of mechanism (B). In a subset of models we identified the expected links to and from the subpolar gyre strength, although only in CESM2 these links are all significant. This subset of models includes those models in which abrupt temperature shifts in the subpolar gyre region have been identified (Swingedouw et al., 2021), indicating that this mechanism likely plays a crucial role in abrupt shifts in the subpolar gyre.
@@
The lagged feedback from the gyre circulation to salinity is not found in any of the models and as a consequence the delayed positive feedback loop is missing. The long memory of the circulation strength can (partly) explain this, as lagged regression (used in Born and Mignot (2012)) can lead to identifying too many significant relationships (McGraw and Barnes, 2018). At a lag of two years we find a weakening effect of the circulation strength on itself in most significant models (shown in Fig. S6), which would act to increase the salinity on a timescale of around 3 years, i.e. a positive feedback loop. This is a shorter timescale than identified by Born and Mignot (2012), which can be due to the gyre-memory biasing their analysis. The gyre-memory can be explained by the transport of waters from outside the Labrador sea, which retains similar properties for multiple years, as seen during the Great Salinity Anomalies (Gelderloos et al., 2012). In CESM2 this lagged link of the gyre strength is found to be significant, providing evidence for both a negative and delayed positive feedback loop being present in this model. Out of the set of models that correctly represent the links to the gyre-strength, also CESM-FV2 and MRI-ESM2-0 show a lagged gyre-strength link, but CESM2-WACCM and NorESM2-LM do not. This does not necessarily mean that the links are not present here, because noise may have reduced their significance.
@@
The link that contradicts most with the bistability mechanism (B) is that from density to the gyre strength. This is one of the links that relies heavily on a good parametrization of heat and salt transport by ocean eddies, and hence it is not surprising that this interaction is not identified clearly. All models considered here have a too coarse resolution to explicitly resolve ocean eddies, and hence we did not find a discernible effect of model resolution on the identification and strength of this link. Another partial explanation is that the barotropic streamfunction, which is used as a measure of the strength of the subpolar gyre, is not the most suitable variable to use for the interaction. The streamfunction is computed by vertical integration, which removes baroclinic effects, whereas it is exactly these baroclinic effects that are relevant for the positive feedback loop (Born and Stocker, 2014). Together, these likely explain the absence of this interaction in most CMIP6 models.
Despite the limitation, it is encouraging to see that the models that best represent mechanism (B) are also the models in which abrupt shifts in the subpolar gyre region are found (Swingedouw et al., 2021). It is important to better understand what sets these models apart, to improve our knowledge about potential bistability in the subpolar gyre and the likelihood of abrupt shifts in this region. Here, exploring the mechanism in eddy-rich models would be a valuable direction for future research to see whether this improves the link between density and the subpolar gyre strength. Furthermore, investigating these mechanisms in an ocean-only model could clarify the relevance of the atmosphere for these interactions, although here one would have to be mindful of the applied atmospheric forcing. Another way to study this would be to include e.g. the North Atlantic Oscillation as a variable forcing the surface variables in mechanism (B) to see the impact this has on the causal links. Furthermore, future research could look at the impact of global warming on the variability mechanism.
@@ Appendix B Convection Mechanisms
Appendix C Subpolar Gyre Variability Mechanism
Link | Lag | Significant Models |
SPGSSS | 1 | ACCESS-CM2, CESM2-FV2, CESM2-WACCM, CESM2, FGOALS-g3, GISS-E2-1-G, GISS-E2-2-G, HadGEM3-GC31-LL |
SSSMLD | 0 | ACCESS-CM2, CESM2-WACCM-FV2, CESM2, CMCC-CM2-SR5, CMCC-ESM2, CanESM5, EC-Earth3-CC, EC-Earth3-Veg-LR, FGOALS-f3-L, GISS-E2-1-G, HadGEM3-GC31-LL, HadGEM3-GC31-MM, MPI-ESM-1-2-HAM, MPI-ESM1-2-LR, MRI-ESM2-0, NorESM2-LM, NorESM2-MM, UKESM1-1-LL |
1 | BCC-CSM2-MR, BCC-ESM1, CESM2-WACCM, CMCC-CM2-SR5, CMCC-ESM2, CanESM5, EC-Earth3-CC, MIROC6, MPI-ESM1-2-LR, MRI-ESM2-0, UKESM1-0-LL | |
MLDSubT | 0 | ACCESS-CM2, BCC-ESM1, CESM2-FV2, CESM2-WACCM-FV2, CESM2, CMCC-CM2-SR5, EC-Earth3-AerChem, EC-Earth3, FGOALS-f3-L, FGOALS-g3, GISS-E2-1-G, GISS-E2-2-G, HadGEM3-GC31-LL, HadGEM3-GC31-MM, IPSL-CM6A-LR, MIROC6, MPI-ESM-1-2-HAM, MPI-ESM1-2-HR, MRI-ESM2-0, NorESM2-LM, NorESM2-MM, UKESM1-0-LL |
1 | CMCC-ESM2, CanESM5, EC-Earth3-CC, EC-Earth3-Veg-LR, EC-Earth3-Veg, IPSL-CM6A-LR, NorESM2-MM | |
SubTRho | 0 | BCC-ESM1, CESM2-FV2, GISS-E2-2-G, HadGEM3-GC31-MM, MPI-ESM1-2-HR, NorESM2-LM, UKESM1-0-LL |
1 | CESM2-FV2, CESM2-WACCM, CESM2, CMCC-CM2-SR5, NorESM2-LM, NorESM2-MM | |
RhoSPG | 0 | EC-Earth3-AerChem, EC-Earth3, GISS-E2-1-G, GISS-E2-2-G, MPI-ESM1-2-HR, MPI-ESM1-2-LR, UKESM1-0-LL |
1 | BCC-CSM2-MR, CESM2-WACCM, CESM2, CMCC-CM2-SR5, CMCC-ESM2, EC-Earth3-CC, FGOALS-f3-L, GISS-E2-1-G, MIROC6, MRI-ESM2-0, NorESM2-LM |
Link | Lag | Significant Models |
SPG | 1 | ACCESS-CM2, BCC-CSM2-MR, BCC-ESM11, CESM2-FV2, CESM2-WACCM, CESM2, CMCC-CM2-SR5, CMCC-ESM2, CanESM5-1, CanESM5, EC-Earth3-AerChem, EC-Earth3-CC, EC-Earth3-Veg-LR, EC-Earth3-Veg, EC-Earth3, FGOALS-f3-L, GISS-E2-1-G, GISS-E2-2-G, HadGEM3-GC31-LL, HadGEM3-GC31-MM, IPSL-CM6A-LR, MIROC6, MPI-ESM-1-2-HAM, MPI-ESM1-2-HR, MPI-ESM1-2-LR, MRI-ESM2-0, NorESM2-LM, NorESM2-MM, UKESM1-0-LL |
2 | BCC-CSM2-MR, BCC-ESM1, CESM2-FV2, CESM2, CanESM5-1, CanESM5, EC-Earth3-AerChem, EC-Earth3-CC, EC-Earth3-Veg-LR, EC-Earth3-Veg, EC-Earth3, GISS-E2-1-G, HadGEM3-GC31-MM, IPSL-CM6A-LR, MIROC6, MPI-ESM1-2-HR, UKESM1-0-LL | |
3 | BCC-CSM2-MR, BCC-ESM1, CESM2-FV2, CESM2, CMCC-CM2-SR5, CanESM5-1, CanESM5, EC-Earth3-AerChem, EC-Earth3-Veg-LR, EC-Earth3-Veg, EC-Earth3, FGOALS-f3-L, FGOALS-g3, HadGEM3-GC31-LL, HadGEM3-GC31-MM, IPSL-CM6A-LR, MIROC6, MPI-ESM1-2-HR, MRI-ESM2-0, UKESM1-0-LL | |
4 | CMCC-ESM2, CanESM5, EC-Earth3-CC, HadGEM3-GC31-MM, IPSL-CM6A-LR, ’MIROC6, MRI-ESM2-0, UKESM1-0-LL | |
5 | CMCC-CM2-SR5, CanESM5-1, CanESM5, EC-Earth3-Veg, HadGEM3-GC31-LL, ’UKESM1-0-LL | |
SSS | 1 | ACCESS-CM2, BCC-CSM2-MR, BCC-ESM1, CESM2-FV2, CESM2-WACCM-FV2, CESM2-WACCM, CESM2, CMCC-CM2-SR5, CMCC-ESM2, CanESM5, FGOALS-g3, ’GISS-E2-1-G, GISS-E2-2-G, HadGEM3-GC31-LL, HadGEM3-GC31-MM, IPSL-CM6A-LR, MIROC6, MPI-ESM-1-2-HAM, MPI-ESM1-2-HR, MPI-ESM1-2-LR, MRI-ESM2-0, NorESM2-LM, NorESM2-MM, UKESM1-0-LL |
MLD | 1 | ACCESS-CM2, BCC-ESM1, CESM2, CMCC-CM2-SR5, CanESM5, EC-Earth3-AerChem, EC-Earth3-CC, EC-Earth3, HadGEM3-GC31-LL, HadGEM3-GC31-MM, IPSL-CM6A-LR, MIROC6, MPI-ESM-1-2-HAM, MPI-ESM1-2-HR, MPI-ESM1-2-LR, MRI-ESM2-0, NorESM2-LM, NorESM2-MM, UKESM1-0-LL |
SubT | 1 | All models |
2 | ACCESS-CM2, CMCC-CM2-SR5, CanESM5-1, CanESM5, EC-Earth3-AerChem, EC-Earth3-CC, EC-Earth3-Veg-LR, EC-Earth3-Veg, EC-Earth3, FGOALS-f3-L, HadGEM3-GC31-LL, IPSL-CM6A-LR, MIROC6, MPI-ESM-1-2-HAM, MPI-ESM1-2-HR, MPI-ESM1-2-LR, MRI-ESM2-0, UKESM1-0-LL | |
Rho | 0 | ACCESS-CM2, BCC-CSM2-MR, BCC-ESM11, CESM2-FV2, CESM2-WACCM-FV2, CESM2-WACCM, CESM2, CMCC-CM2-SR5, CanESM5-1, CanESM5, EC-Earth3-AerChem, EC-Earth3-CC, EC-Earth3-Veg-LR, EC-Earth3-Veg, EC-Earth3, FGOALS-f3-L, FGOALS-g3, GISS-E2-1-G, GISS-E2-2-G, HadGEM3-GC31-LL, HadGEM3-GC31-MM, IPSL-CM6A-LR, MIROC6, MPI-ESM-1-2-HAM, MPI-ESM1-2-HR, MPI-ESM1-2-LR, MRI-ESM2-0, NorESM2-LM, NorESM2-MM, UKESM1-0-LL, UKESM1-1-LL |